OpenStudy (anonymous):

@pooja195 @TheSmartOne may you help me check my homework?

OpenStudy (anonymous):

If f(x) is the total cost, in dollars, of x candies, which of the following statements best describes the meaning of f(2) = 6? The total cost of 2 candies is $6.00. The total cost of 6 candies is $2.00. The total cost of 2 candies is $3.00. The total cost of 3 candies is $2.00.

OpenStudy (anonymous):

a?

pooja195 (pooja195):

\(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\sf\color{blue}{~~~~Welcome~to~OpenStudy!~\ddot\smile}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\)

pooja195 (pooja195):

yes A is right

OpenStudy (anonymous):

The first five terms of a sequence are shown below: 6, 9, 12, 15, 18 If the nth term of this sequence is represented by f(n), which of the following functions best represents this sequence? f(n) = 6n - 3; n ≥ 1 f(n) = 6n + 3; n ≥ 1 f(n) = 3n + 3; n ≥ 1 f(n) = 3n - 3; n ≥ 1

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

is this a test?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

review packet @pooja195

pooja195 (pooja195):

@TheUnSmartOne please help

pooja195 (pooja195):

@TheSmartOne

OpenStudy (anonymous):

@TheSmartOne